OpenStudy (anonymous):

A triangle has a leg that measures 6 and a hypotenuse of 12. what measure would form a right triangle?

OpenStudy (anonymous):

\[a^2+b^2=c^2\] c=hypotenuse a and b are the rest of the sides.

OpenStudy (anonymous):

c=12 a=6 b is the side you are looking for. Plug in the values and solve for b

OpenStudy (anonymous):

h=12 base=6 height=? applying pythagorus-\[12^{2}=6^{2}+height^{2}\] \[12^{2}-6^{2}=height^{2}\] \[144-36=height^{2}\] \[108=height^{2}\] \[\sqrt{108}=height\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

How is this possible you may ask. This is pythagoras theorem. If you want more information you can just google pythagoras theorem